OpenStudy (anonymous):

5(x-5) over 2 +7=16 Does anyone know how to do this?

OpenStudy (anonymous):

open parenthesis: (5x-25)/2 +7 =16 multiply everything by 2 and get (5x-25) +14 = 32 can you proceed from there?

OpenStudy (anonymous):

8.6

OpenStudy (anonymous):

what do you do next?

OpenStudy (anonymous):

put like terms together 5x = 32-14+25 5x = 43 divide all through by 5 and get x x=??

OpenStudy (anonymous):

oh okay thanks.

OpenStudy (anonymous):

Don't give ANSWERS!!!

OpenStudy (compassionate):

TO SOLVE AN EQUATION WITH fractions, we transform it into an equation without fractions -- which we know how to solve. The technique is called clearing of fractions. Multiply both sides of the equation -- every term -- by the LCM of denominators. Every denominator will then cancel. We will then have an equation without fractions.